# Time-dependent Predictive Accuracy Metrics in the Context of Interval Censoring and Competing Risks

### v1

@@ 2.3 Sensitivity, specificity and AUC
$\mathcal{W^{\prime}}^{\textsc{m}}_{i}=\begin{cases}\Pi^{\textsc{prg}}_{i}(\infty\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{cen}}_{i}>t,T^{\textsc{cen}}_{i}<t+\Delta t\quad\scriptstyle{(1c)},\\ \Pi^{\textsc{prg}}_{i}(\infty\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{cen}}_{i}<t+\Delta t\quad\scriptstyle{(3c)},\\ \Pi^{\textsc{prg}}_{i}(\infty\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{prg-}}_{i}<t+\Delta t,T^{\textsc{cen}}_{i}>t+\Delta t\quad\scriptstyle{(2c)},\\ \Pi^{\textsc{prg}}_{i}(T^{\textsc{prg+}}_{i}\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{prg-}}_{i}<t+\Delta t,T^{\textsc{prg+}}_{i}>t+\Delta t\quad\scriptstyle{(2a)},\\ \Pi^{\textsc{prg}}_{i}(T^{\textsc{trt}}_{i}\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{prg-}}_{i}<t+\Delta t,T^{\textsc{trt}}_{i}<t+\Delta t\quad\scriptstyle{(2b)},\\ \Pi^{\textsc{prg}}_{i}(\infty\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{cen}}_{i}>t+\Delta t\quad\scriptstyle{(5c)},\\ \Pi^{\textsc{prg}}_{i}(T^{\textsc{prg+}}_{i}\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{prg+}}_{i}>t+\Delta t\quad\scriptstyle{(5a)},\\ \Pi^{\textsc{prg}}_{i}(T^{\textsc{trt}}_{i}\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{trt}}_{i}>t+\Delta t\quad\scriptstyle{(5b)},\\ 1&\text{if }T^{\textsc{prg-}}_{i}>t+\Delta t\quad\scriptstyle{(4a,4b,4c)},\\ 0&\text{otherwise}\quad\scriptstyle{(1a,1b,3a,3b,\text{ and the greyed-out scenarios})}.\end{cases}$
The model-based weights for being a control involve predictions of the progression-specific risk at infinity. Often, this is approximated by replacing infinity with a large enough number. Depending on the model or algorithm, however, this extrapolation can be very unstable. For example, the baseline hazard function in the proportional hazard submodel of a joint model is often parametrized using the P-splines,(Rizopoulos and Ghosh, 2011) which might not approximate the hazard well for times outside the time range in the training set (see Section 3). One alternative approach is to use the risk prediction for the last time point in the training set carried forward. However, that is a rather strict and typically very unrealistic assumption since the risk likely increases over time. To overcome this problem, we rely on a less strict assumption, namely, that the ratio between the risk of the primary and competing event remains stable outside the time range covered by the data. Since the sum of the cumulative risks for both events goes to one as time goes to infinity, the assumption of a constant ratio between the risks is sufficient to obtain an estimate for the progression-specific risk at infinity.
@@
Figure 2: Illustration of the model-based weights for sensitivity (a) and specificity (b) in calculating AUC and Brier score. The shaded bars represent the time with which the patients contribute to being a case (a) or a control (b). Grayed-out patient groups are not considered in the calculation.
@@
$\widehat{\text{sen}}^{\textsc{prg}}_{\textsc{ipcw}}(t,\Delta t,c)$ $=\frac{\sum_{i:T^{\textsc{prg-}}_{i}\geq t,T^{\textsc{prg+}}_{i}<t+\Delta t}I\{\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)\geq c\}\times\mathcal{W}^{\textsc{ipcw}}_{i}}{\mathcal{W}^{\textsc{ipcw}}_{i}},$
where $\mathcal{W}^{\textsc{ipcw}}_{i}=\{G(T^{\textsc{prg+}}_{i}\mid t)\}^{-1}$, and $G(\cdot)$ is the conditional cumulative probability of being censoring-free. Furthermore, to estimate the IPCW-based progression-specific specificity, the subset of the patients known to be event-free at $t+\Delta t$, i.e., the absolute controls (4a, 4b and 4c in Figure 1), are used
$\widehat{\text{spe}}^{\textsc{prg}}_{\textsc{ipcw}}(t,\Delta t,c)$ $=\frac{\sum_{i:T^{\textsc{prg-}}_{i}>t+\Delta t}I\{\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)<c\}\times\mathcal{W^{\prime}}^{\textsc{ipcw}}_{i}}{\mathcal{W^{\prime}}^{\textsc{ipcw}}_{i}}.$
The corresponding weights are estimated using the KM estimator for being censoring-free at time $t+\Delta t$, to represent the patients censored during the interval of interest. Thus the weights are the inverse of the conditional probability of being censoring-free at the end of the interval of interest, $\mathcal{W^{\prime}}^{\textsc{ipcw}}_{i}=\{G(t+\Delta t\mid t)\}^{-1}$.
@@ 2.4 Brier score
$\text{BS}^{\textsc{prg}}(t+\Delta t,t)=E\left[\left\{I(T^{\textsc{prg*}}_{i}<t+\Delta t)-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)\right\}^{2}\mid T^{(2)}_{i}\geq t,T^{\textsc{prg*}}_{i}<T^{\textsc{trt*}}_{i}\right],$
@@
where $n_{t}$ is the number of patients who have not progressed, been treated, or censored until time $t$.
@@ 3 Application to the ICJM
To illustrate the practical application of these methods, we analyze a joint model fitted on data from the Canary Prostate Active Surveillance Study. This study focuses on the progression of prostate cancer to a stage necessitating intervention, with progression detected through interval-censored biopsies. Additionally, the initiation of early treatment, which occurs before the cancer reaches a critical stage, is considered a competing risk. The patients were examined with biopsies according to the PASS schedules (at 12, 24 months, and biennially afterwards) to detect the primary event of interest, cancer progression, resulting in interval censoring. Around 10% of the patients left AS for treatment before cancer progression was detected, constituting the competing event.(Newcomb et al., 2016) In addition, the repeatedly measured biomarker, PSA, is a time-varying predictor for the events of interest. We developed the ICJM 1 in Yang et al. (2024a) from the Canary PASS data. The repeatedly measured PSA levels were modelled in the longitudinal submodel of the ICJM, a mixed-effects model,
$\log_{2}\{\mbox{PSA}_{i}(t)+1\}=\beta_{0}+u_{0i}+\displaystyle{\sum}^{3}_{p=1}(\beta_{p}+u_{pi})\mathcal{C}^{(p)}_{i}(t)+\beta_{4}(\mbox{Age}_{i}-62)+\epsilon_{i}(t),\quad\boldsymbol{u}_{i}\sim N(\mathbf{0},\boldsymbol{\Omega}),$
where $\mathcal{C}(\cdot)$ is the design matrix for the natural cubic splines, $\mbox{Age}_{i}$ refers to the patient’s age at the start of active surveillance, and $\boldsymbol{\beta}=(\beta_{0},\dots,\beta_{4})^{\top}$ is a vector of corresponding regression coefficients. The error terms $\epsilon_{i}(t)$ were assumed to follow a Student’s t-distribution with three degrees of freedom, and the random effects, $\boldsymbol{u}_{i}=\{u_{0i},u_{1i},u_{2i},u_{3i}\}^{\top}$, were assumed to be normally distributed with a mean of zero and an unstructured variance-covariance matrix $\boldsymbol{\Omega}$. The expected values of $\log_{2}\{\mbox{PSA}_{i}(t)+1\}$, which we denote by $m_{i}(t)$, were then incorporated into a survival submodel using two functional forms:
@@
where $\mathcal{G}_{a}(t,\boldsymbol{\xi})$ is the $a$-th basis function of a B-splines with knots $\xi_{1},\dots,\xi_{A}$. The number of knots was chosen to be 11, and the penalized coefficients for the basis function are denoted by $\boldsymbol{\gamma}_{k,h_{0}}$. The ICJM was estimated in the Bayesian framework. We evaluated the predictive performance of the ICJM for the interval of interest, $(t,t+\Delta t]=(1,4]$, i.e., using covariate information up until year 1. The evaluation was performed internally on the Canary PASS data using 5-fold cross-validation. A total of 812 (out of 833) subjects at risk in year one were included in the evaluation. During the interval of interest, 50 subjects started the competing event, early treatment. The model based approach resulted in an AUC of 0.63 while the AUC estimated using the IPCW approach was higher, namely 0.65. Also in the calculation of the Brier score, the model based approach estimated a less favorable performance than the IPCW approach, namely 0.10 versus 0.07. The model’s estimated EPCE was -0.71.
@@ 4.1 Simulation setting
As the results of the application in Section 3 show,, the model-based and IPCW approach can result in relatively different estimates of AUC and Brier score. To investigate the performance of the two methods in a more controlled setting, we conducted two simulation studies. In both studies, the accuracy metrics were calculated using the longitudinal information (i.e., PSA) until year one and evaluating the performance up until year four, i.e., the interval of interest being $[1,4)$. 200 datasets were simulated based on the ICJM described in Section 3 and the PASS biopsy schedule (Newcomb et al., 2016), each with 300 subjects.
@@ 4.2 Comparison between model specification
Since the model-based approach re-uses the prediction model for its evaluation, model misspecification might lead to overestimation of the model’s predictive performance. Therefore, in this simulation, we aim to explore the influence of model misspecification on the accuracy metrics from both approaches. Three models, including the correctly-specified model and two misspecified models (i.e., ignoring one baseline covariate in the survival submodel and wrongly assuming a linear evolution for the PSA trajectory), were fitted on each dataset. Each fitted model was then evaluated on another simulated dataset. As a reference, we also calculated AUC, Brier score, and EPCE, based on the true event times, i.e., for the ideal scenario where there is no censoring, and it is, thus, known whether a patient is a case or a control.
The main results from the simulation study are shown in Figure 3 and Table 1. Comparison of the AUC between the three models for the reference scenario without censoring (Figure 3(a)) shows that misspecification of the effect of time on PSA as linear had no relevant effect on the model’s performance whereas omitting the baseline covariate worsened the discriminative ability considerably. The IPCW estimates of the AUC generally had greater variability whereas the median estimates across all three models were closer to the reference AUC than the median of the model-based estimates. The model-based approach tended to overestimate the AUC. However, the variability of the model-based AUC across simulations was not inflated. The RMSE of the AUC from the IPCW approach was stable across all three models (approximately 0.054) while the RMSE of the model-based AUC was lower for the correct model and the model with mis-specified PSA slope (0.047 and 0.043, respecitvely) but larger for the model omitting the baseline covariate (0.058). The estimates of the Brier score were not much affected by model misspecification. Both of the approaches underestimated the Brier score. Analogous to the AUC, the model-based Brier scores for all three models had less variability but were, on average, closer to the reference Brier score than the one from the IPCW approach. Consequently, the RMSE of the model-based Brier scores were lower than those of the IPCW Brier scores for all three models. The model-based estimates of the EPCE for the three models were comparable to the reference EPCE (see Figure 3(c)). The reference EPCE also did not vary much among different models.
@@
Figure 3: The model-based and IPCW estimates of the AUC (a) and Brier score (b), and the model-based estimates of the EPCE (c) from the correctly-specified model, the model with a linear PSA trajectory, and the model ignoring the baseline covariate (PSA density), compared to the corresponding ideal scenario when there is no censoring.
@@
Model-based IPCW
Correctly-specified 0.047 0.053
Linear 0.043 0.054
No baseline covariate 0.058 0.054
@@
Model-based IPCW
Correctly-specified 0.079 0.086
Linear 0.080 0.086
No baseline covariate 0.080 0.088
Table 1: The root mean square error (RMSE) of the model-based and the IPCW estimations of the AUC (a) and Brier score (b) from the correctly-specified model, the model with a PSA trajectory, and the model ignoring the baseline covariate (PSA density), to the corresponding reference metrics without censoring.

### v2

@@ 2.3 Sensitivity, specificity and AUC
$\mathcal{W^{\prime}}^{\textsc{m}}_{i}=\begin{cases}\mathcal{S}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{cen}}_{i}>t,T^{\textsc{cen}}_{i}<t+\Delta t\quad\scriptstyle{(1c)},\\ \frac{\Pi^{\textsc{prg}}_{i}(T^{\textsc{prg+}}_{i}\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})}{\Pi^{\textsc{prg}}_{i}(T^{\textsc{prg+}}_{i}\mid T^{\textsc{prg-}}_{i})}&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{prg-}}_{i}<t+\Delta t,T^{\textsc{prg+}}_{i}>t+\Delta t\quad\scriptstyle{(2a)},\\ 1-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{prg-}}_{i}<t+\Delta t,T^{\textsc{trt}}_{i}>t+\Delta t\quad\scriptstyle{(2b)},\\ 1-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{prg-}}_{i}<t+\Delta t,T^{\textsc{cen}}_{i}>t+\Delta t\quad\scriptstyle{(2c)},\\ \mathcal{S}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{cen}}_{i}<t+\Delta t\quad\scriptstyle{(3c)},\\ 1&\text{if }T^{\textsc{prg-}}_{i}>t+\Delta t\quad\scriptstyle{(4a,4b,4c)},\\ \frac{\Pi^{\textsc{prg}}_{i}(T^{\textsc{prg+}}_{i}\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})}{\Pi^{\textsc{prg}}_{i}(T^{\textsc{prg+}}_{i}\mid T^{\textsc{prg-}}_{i})}&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{prg+}}_{i}>t+\Delta t\quad\scriptstyle{(5a)},\\ 1-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{trt}}_{i}>t+\Delta t\quad\scriptstyle{(5b)},\\ 1-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{cen}}_{i}>t+\Delta t\quad\scriptstyle{(5c)},\\ 0&\text{otherwise}\quad\scriptstyle{(1a,1b,3a,3b,\text{ and the grayed-out scenarios})}.\end{cases}$
The model-based weights of being controls for the patients who had progression detected and can potentially be a control (i.e., patient group 2a and 5a) are constructed analogously to their counterpart of being cases: the differences of progression-specific risks between the progression detection time ($T^{\textsc{prg+}}_{i}$) and $t+\Delta t$ scaled relative to the risk differences between the progression detection time ($T^{\textsc{prg+}}_{i}$) and the last negative biopsy time ($T^{\textsc{prg-}}_{i}$. For patient groups 1c and 3c who were censored before the end of the interval of interest (i.e., $t+\Delta t$), the model-based weights for being controls are the probability of the patient not having either of the events before $t+\Delta t$ conditional on the fact that neither event happened until the last negative biopsy, i.e., the conditional overall survival function $\mathcal{S}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})$. For patients who were censored or treated after $t+\Delta t$ (i.e., patient groups 2b, 2c, 5b and 5c), it is known that they did not initiate treatment before $t+\Delta t$. Therefore, their weight of being a control is calculated as one minus the conditional progression-specific risk at $t+\Delta t$.
@@
Figure 2: Illustration of the model-based weights for sensitivity (a) and specificity (b) in calculating AUC and Brier score. The shaded bars represent the time with which the patients contribute to being a case (a) or a control (b). The lightgray shades are the weights using the predicted progression-specific risks. The dark shades indicate using the rescaled progression-specific risks. The crosshatches indicate using the overall survival function. Grayed-out patient groups are not considered in the calculation.
@@
$\widehat{\text{sen}}^{\textsc{prg}}_{\textsc{ipcw}}(t,\Delta t,c)$ $=\frac{\sum_{i:T^{\textsc{prg-}}_{i}\geq t,T^{\textsc{prg+}}_{i}<t+\Delta t}I\{\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)\geq c\}\times\mathcal{W}^{\textsc{ipcw}}_{i}}{\sum_{i:T^{\textsc{prg-}}_{i}\geq t,T^{\textsc{prg+}}_{i}<t+\Delta t}\mathcal{W}^{\textsc{ipcw}}_{i}},$
where $\mathcal{W}^{\textsc{ipcw}}_{i}=\{G(T^{\textsc{prg+}}_{i}\mid t)\}^{-1}$, and $G(s\mid t)$ is the conditional probability of a patient being censoring-free at $s$ given being uncensored at $t$. Furthermore, to estimate the IPCW-based specificity, the subset of the patients known to be event-free at $t+\Delta t$, i.e., the absolute controls (4a, 4b and 4c in Figure 1), is used
$\widehat{\text{spe}}_{\textsc{ipcw}}(t,\Delta t,c)$ $=\frac{\sum_{i:T^{\textsc{prg-}}_{i}>t+\Delta t}I\{\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)<c\}\times\mathcal{W^{\prime}}^{\textsc{ipcw}}_{i}}{\sum_{i:T^{\textsc{prg-}}_{i}>t+\Delta t}\mathcal{W^{\prime}}^{\textsc{ipcw}}_{i}}.$ (3)
The corresponding weights are estimated using the KM estimator for being censoring-free at time $t+\Delta t$, to represent the patients censored during the interval of interest. Thus the weights are the inverse of the conditional probability of being censoring-free at the end of the interval of interest, $\mathcal{W^{\prime}}^{\textsc{ipcw}}_{i}=\{G(t+\Delta t\mid t)\}^{-1}$. As these weights are uniform for all the controls and will unsurprisingly cancel out, (3) can be simplified to
$\widehat{\text{spe}}_{\textsc{ipcw}}(t,\Delta t,c)$ $=\frac{\sum_{i:T^{\textsc{prg-}}_{i}>t+\Delta t}I\{\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)<c\}}{\sum_{i:T^{\textsc{prg-}}_{i}>t+\Delta t}1}.$
@@ 2.4 Brier score
$\text{BS}^{\textsc{prg}}(t+\Delta t,t)=E\left[\left\{I(T^{\textsc{prg*}}_{i}<t+\Delta t)-\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)\right\}^{2}\mid T^{(2)}_{i}\geq t\right],$
@@
where $n_{t}$ is the number of patients who have not been detected with progression, been treated, or censored until time $t$.
@@ 3 Application to the ICJM
To illustrate the practical application of these methods, we analyze a joint model that can handle interval censoring and competing risks fitted on data from the Canary Prostate Active Surveillance Study. This study focuses on the progression of prostate cancer to a stage necessitating intervention, with progression detected through interval-censored biopsies. Additionally, the initiation of early treatment, which occurs before the cancer reaches a critical stage, is considered a competing risk. The patients were examined with biopsies according to the PASS schedules (at 12, 24 months, and biennially afterwards) to detect the primary event of interest, cancer progression, resulting in interval censoring. Around 10% of the patients left AS for treatment before cancer progression was detected, constituting the competing event.(Newcomb et al., 2016) In addition, the repeatedly measured biomarker, PSA, is a time-varying predictor for the events of interest. We developed the ICJM 1 in Yang et al. (2025) from the Canary PASS data (more details can be found in Web Appendix 2). The repeatedly measured PSA levels, for patient $j$ in the training set, were modelled in the longitudinal submodel of the ICJM, a mixed-effects model,
$\log_{2}\{\mbox{PSA}_{j}(t)+1\}=\beta_{0}+u_{0j}+\displaystyle{\sum}^{3}_{p=1}(\beta_{p}+u_{pj})\mathcal{C}^{(p)}_{j}(t)+\beta_{4}(\mbox{Age}_{j}-62)+\epsilon_{j}(t),\quad\boldsymbol{u}_{j}\sim N(\mathbf{0},\boldsymbol{\Omega}),$
where $\mathcal{C}(\cdot)$ is the design matrix for the natural cubic splines, $\mbox{Age}_{j}$ refers to the patient’s age at the start of active surveillance, and $\boldsymbol{\beta}=(\beta_{0},\dots,\beta_{4})^{\top}$ is a vector of corresponding regression coefficients. The error terms $\epsilon_{j}(t)$ were assumed to follow a Student’s t-distribution with three degrees of freedom, and the random effects, $\boldsymbol{u}_{j}=\{u_{0i},u_{1i},u_{2i},u_{3i}\}^{\top}$, were assumed to be normally distributed with a mean of zero and an unstructured variance-covariance matrix $\boldsymbol{\Omega}$. The expected values of $\log_{2}\{\mbox{PSA}_{j}(t)+1\}$, which we denote by $m_{j}(t)$, were then incorporated into a survival submodel using two functional forms:
@@
where $\mathcal{G}_{a}(t,\boldsymbol{\xi})$ is the $a$-th basis function of a B-splines with knots $\xi_{1},\dots,\xi_{A}$. The number of knots was chosen to be 11, and the penalized coefficients for the basis function are denoted by $\boldsymbol{\gamma}_{k,h_{0}}$. The ICJM was estimated in the Bayesian framework. We evaluated the predictive performance of the ICJM for the interval of interest, $(t,t+\Delta t]=(1,4]$, i.e., using covariate information up until year 1. The evaluation was performed internally on the Canary PASS data using 5-fold cross-validation. A total of 812 (out of 833) subjects at risk in year one were included in the evaluation. During the interval of interest, 50 subjects started the competing event, early treatment. The model-based approach resulted in an AUC of 0.64 while the AUC estimated using the IPCW approach was higher, namely 0.68. Also in the calculation of the Brier score, the model-based approach estimated a less favorable performance than the IPCW approach, namely 0.17 versus 0.07. The model’s estimated EPCE was -0.72.
@@ 4.1 Simulation setting
As the results of the application in Section 3 show, the model-based and IPCW approach can result in relatively different estimates of AUC and Brier score. To investigate the performance of the two methods in a more controlled setting, we conducted two simulation studies. In both studies, the accuracy metrics were calculated using the longitudinal information (i.e., PSA) until year one in the test set and evaluating the performance up until year four, i.e., the interval of interest being $[1,4)$. 200 datasets were simulated based on the ICJM handling interval censoring of biopsies and competing risk of early treatment described in Section 3 and the PASS biopsy schedule (Newcomb et al., 2016), each with 300 subjects. In the 200 datasets, there are on average 22% patients who were observed with cancer progression, 9% who initiated early treatment.
@@ 4.2 Comparison between model specification
Since the model-based approach re-uses the prediction model for its evaluation, model misspecification might lead to overestimation of the model’s predictive performance. Therefore, in this simulation, we aim to explore the influence of model misspecification on the accuracy metrics from both approaches. Three models, including the correctly-specified model and two misspecified models (i.e., ignoring one baseline covariate in the survival submodel and wrongly assuming a linear evolution for the PSA trajectory), were fitted on each dataset (as the training set). Each fitted model was then evaluated on another simulated dataset (as the test set in which we utilize the PSA information until year one for predictions). As a reference, we also calculated AUC, Brier score, and EPCE, based on the true event times, i.e., for the ideal scenario where there is no censoring, and it is, thus, known whether a patient is a case or a control. It is noted that the reference metrics were calculated using the same risk estimates from the corresponding model when compared to the model-based and IPCW approach. Furthermore, we also presented a naive calculation of AUC and Brier score, i.e., ignoring the interval censoring essence of cancer progression and using the observed progression time to determine a case or control.
The main results from the simulation study are shown in Figure 3 and Table 1. Comparison of the AUC between the three models for the ideal scenario without censoring (Figure 3(a)) shows that misspecification of the effect of time on PSA as linear had no relevant effect on the model’s performance whereas omitting the baseline covariate worsened the discriminative ability considerably. Simply ignoring interval censoring resulted in, on average, relatively small differences in the AUC compared to the ideal scenario in which the exact event times are known. However, the variability in this approach was larger than in the model-based approach. The naive approach performed worse than the model-based approach but was comparable to the IPCW approach in estimating the Brier score for all three models whether misspecified or not. The IPCW estimates of the AUC generally had greater variability whereas the median estimates across all three models were closer to the reference AUC than the median of the model-based estimates. The model-based approach tended to overestimate the AUC. Misspecification of the model by omitting the baseline covariate resulted in a larger RMSE for the model-based approach. However, the variability of the model-based AUC across simulations was not inflated. Analogous to the AUC, the model-based Brier scores for all three models had less variability but were, on average, closer to the reference Brier score than the one from the IPCW approach. Only the model-based approach generated estimates of the Brier score that were similar to the ideal scenario for all three models. The model-based approach also had the lowest RMSE in estimating both the AUC and Brier scores for all three models, approximately 38% better for the AUC and 83% better for the Brier score compared to the IPCW approach, and approximately 24% better for the AUC and 82% better for the Brier score compared to the naive approach. The model-based estimates of the EPCE for the three models were comparable to the reference EPCE (see Figure 3(c)). The reference EPCE also did not vary much among different models.
@@
Figure 3: The model-based, IPCW and naive (ignoring interval censoring) estimates of the AUC (a) and Brier score (b), and the model-based estimates of the EPCE (c) from the correctly-specified model, the model with a linear PSA trajectory, and the model ignoring the baseline covariate (PSA density), compared to the corresponding ideal scenario when there is no censoring.
@@
Model-based IPCW Naive
Correctly-specified 0.036 0.056 0.045
Linear 0.029 0.054 0.043
No baseline covariate 0.040 0.058 0.051
@@
Model-based IPCW Naive
Correctly-specified 0.015 0.087 0.083
Linear 0.015 0.086 0.082
No baseline covariate 0.015 0.088 0.085
Table 1: The root mean square error (RMSE) of the model-based, IPCW and naive (ignoring interval censoring) estimates of the AUC (a) and Brier score (b) from the correctly-specified model, the model with a PSA trajectory, and the model ignoring the baseline covariate (PSA density), to the corresponding reference metrics without censoring.